OpenStudy (anonymous):

verify (sinx)/(cscx) -1 = sinxcot(-x)cos(-x)

OpenStudy (anonymous):

did you do any work ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

pls show your steps

OpenStudy (anonymous):

(sinx)/(cscx) - (cscx)/(cscx)

OpenStudy (anonymous):

(sinX-cscx)/(cscx)

OpenStudy (anonymous):

= sin^2 x -1= -cos^2x

OpenStudy (anonymous):

now simplify right side of equation

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

sin^x + cos^x=1, so sin^x-1= - cos^2 x

OpenStudy (anonymous):

how did u get cosx

OpenStudy (anonymous):

or the 1

OpenStudy (anonymous):

i just gave the identity

OpenStudy (anonymous):

i know but we can't use that identity

OpenStudy (anonymous):

who said that

OpenStudy (anonymous):

all necesary steps have been show, if a^2+b^2=1 , then a^2-1= -b^2

OpenStudy (anonymous):

what i meant was where did u use that identity

OpenStudy (anonymous):

well you had sin x / cosec x - cosec x / cosec x= sin x * sin x -1, as cosec x = 1/sin x

OpenStudy (anonymous):

i think this is not true

OpenStudy (anonymous):

it is absolutely true, but for the lack of efforts shown, anyways work from right side of equation and you will have -cosx * -cos x = -cos^2 x

OpenStudy (anonymous):

on the right side i got (six -cscx)/cscx then wht do i do

OpenStudy (anonymous):

sin x / cot (-x) = -cos x cot(-x)=-cot x BUT cos(-x)= cos(x)

OpenStudy (anonymous):

sorry i got the right side but not the left side

OpenStudy (anonymous):

how can i get (sinx-cscx)/cscx to equal -cos^2 x

OpenStudy (anonymous):